研究集会「リー群の表現論」

不変微分作用素，指標， Orbit Method をめぐって

Representation Theory of Lie Groups

— In honor of Professor Takeshi HIRAI —

Dates: February 18 (Monday) – February 22 (Friday), 2002 Location: Research Institute for Mathematical Sciences (Room 420) Kyoto University, Sakyo, Kyoto 606-8502, Japan

Organizers: K. Nishiyama, H. Yamashita, M. Kashiwara

Program:

time  Feb.18(Mon) Feb.19(Tue) Feb.20(Wed) Feb.21(Thu) Feb.22(Fri)
9:30 – 10:30  Fujiwara  Takase  {Ochiai \atop (9:20–10:20)}  Brion  Nishiyama
10:45 – 11:45  Nomura  Schiffmann {Kato \atop (10:30–11:30)} Kawanaka  Knop
lunch  lunch  {Uzawa \atop (11:45–12:45)}  lunch  lunch
13:30 – 14:30  Itoh  Torasso  No lecture  Ohta  Yamashita
14:45 – 15:45  Oshima  Aoki-Kato  No lecture  Sekiguchi  Loke
16:00 – 17:00  Duflo  Hirai  No lecture  Levasseur  Matumoto

{Banquet \atop (18:00 – )}

February 18 (Mon)

9:25 – 9:30 Opening ceremony

9:30 – 10:30 Hidenori Fujiwara (Kinki University)

“Analysis on nilpotent homogeneous spaces; reciprocity and differential operators”

10:45 – 11:45 Takaaki Nomura (Kyoto University)

“Symmetry characterizations for homogeneous Siegel domains related to Laplace-Beltrami operators ”

13:30 – 14:30 Minoru Itoh (Kyoto University)

“Capelli identities for the dual pair (O_{M}, Sp_{N})”

14:45 – 15:45 Toshio Oshima (University of Tokyo)

“A quantization of coadjoint orbits and differential equations attached to generalized flag manifolds”

16:00 – 17:00 Michel Duflo (Université Paris VII)

“On Poincaré-Birkhoff-Witt theorem and CDYBE equation”

February 19 (Tue)

9:30 – 10:30 Kouichi Takase (Miyagi University of Education)

“On Fourier transform of spherical trace function”

10:45 – 11:45 Gérard Schiffmann (Université de Strasbourg, Professor emeritus) “Invariant distributions, involutions and multiplicity one theorems”

13:30 – 14:30 Pierre Torasso (Université de Poitiers)

“The character formula for almost algebraic groups”

14:45 – 15:45 Shigeru Aoki (Takushoku Univ.) and Suehiro Kato (Kitasato Univ.) “Matching conditions for invariant eigendistributions on semisimple symmetric spaces ”

16:00 – 17:00 Takeshi Hirai (Kyoto University, Professor emeritus)

“On unitary representations of infinite classical Weyl groups”

18:00 – Banquet (Conference dinner) at Kyôdai-kaikan

February 20 (Wed)

9:20 – 10:20 Hiroyuki Ochiai (Tokyo Institute of Technology)

“Eigenspace representations of symmetric spaces of exceptional type”

10:30 – 11:30 Syu Kato (University of Tokyo)

“On the higher cohomology group of group completion”

11:45 – 12:45 Tohru Uzawa (Rikkyo University)

“Symmetric varieties”

February 21 (Thu)

9:30 – 10:30 Michel Brion (Université Joseph Fourier)

“A geometric approach to standard monomial theory”

10:45 – 11:45 Noriaki Kawanaka (Osaka University)

“Sato-Welter game and its generalizations

— 2-person games for representation theorists”

13:30 – 14:30 Takuya Ohta (Tokyo Denki University)

“Nilpotent orbits of ℤ_{4}-graded Lie algebra and geometry of moment maps associated to the dual pair (U(p, q), U(r, s)), (O(p, q), Sp(2n, ℝ)) and (Sp(p, q), O^{∗}(2n)) ”

14:45 – 15:45 Jiro Sekiguchi (Tokyo University of Agriculture and Technology) “Recent progress on nilpotent orbits of complex symmetric pairs”

16:00 – 17:00 Thierry Levasseur (Université de Brest)

“Differential operators on highest weight varieties”

February 22 (Fri)

9:30 – 10:30 Kyo Nishiyama (Kyoto University)

“Theta lifting of nilpotent orbits and multiplicity-free action”

10:45 – 11:45 Friedrich Knop (Rutgers University)

“Harmonic analysis on multiplicity free spaces”

13:30 – 14:30 Hiroshi Yamashita (Hokkaido University)

“Isotropy representation attached to Harish-Chandra module with irreducible associated variety”

14:45 – 15:45 Hung-Yean Loke (National University of Singapore/ Kyoto Univ.) “Howe quotients of unitary lowest weight modules”

16:00 – 17:00 Hisayosi Matumoto (University of Tokyo)

“On the socle of degenerate principal series representation of U(m, n)”

The latest information will be found in the conference web-page

http://w3rep.math.h.kyoto-u.ac.jp/meeting/wshirai2002.html

For further information, please contact

Kyo Nishiyama Faculty of Integrated Human Studies Kyoto University Sakyo, Kyoto 606-8501, Japan kyo@math.h.kyoto-u.ac.jp or  Hiroshi Yamashita Graduate School of Science Hokkaido University Sapporo 060-0810, Japan yamasita@math.sci.hokudai.ac.jp